OpenStudy (anonymous):

Which of the following will give you the incorrect slope? A: y2-y1 B: y2-y1 C: y1-y2 D: Rise --------- -------- -------- -------- x2-x1 x1-x2 x1-x2 Run

OpenStudy (anonymous):

Anyone? :(

OpenStudy (anonymous):

It is the second one because the order of subtraction is reversed in the denominator. The subscripts have to line up. Selection 4 is fine because that is the same as the difference in y over the difference in x.

OpenStudy (anonymous):

Notice how in A and C, the order of the subscripts is the same, which makes those answers good.

OpenStudy (anonymous):

Thank you soooo much!

OpenStudy (anonymous):

You're welcome!

OpenStudy (anonymous):

Can you help me with one more?

OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (−6, −6) and (−3, 1) A:−7/3 B: 3/7 C:7/3 D: -3/7

OpenStudy (anonymous):

You would need to use the slope formula:\[\frac{ y _{1} - y _{2} }{ x _{1} - x _{2} } = m\]where "m" is the slope and \[(x _{1}, y _{1})\]and\[(x _{2}, y _{2})\]are your given points. So, x1= -6, y1 = -6, x2 = -3, and y2 = 1. Just make those substitutions and you will be able to get your answer.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

But is that clear enough to follow? I don't just want to leave you with a formula. There is a good methodology here, but can you make the substitutions? Is this helping you?

OpenStudy (anonymous):

Haha, i'm kinda confused..

OpenStudy (anonymous):

Nvm i got it!

OpenStudy (anonymous):

Excellent! Good job! I suggest memorizing this formula because it is one of the most used and useful in math at any level.

OpenStudy (anonymous):

Wait.. Would the answer be -7/3?

OpenStudy (anonymous):

Actually, it is 7/3, what you had but without the minus sign. You're on the right track, but be careful about your signs.

OpenStudy (anonymous):

Thank you! i really appreciate your help :)

OpenStudy (anonymous):

\[\frac{ -6 - 1 }{ -6 - (-3) } = \frac{ -6 - 1 }{ -6 + 3 } = \frac{ -7 }{ -3 } = \frac{ 7 }{ 3 }\]

OpenStudy (anonymous):

You're very welcome, @ShayTaylorO But take a look at how I resolved the sign.

OpenStudy (anonymous):

I see now! Thanks!!! :D

OpenStudy (anonymous):

Cool! Very nice working with you and I wish you luck in all of your studies! @ShayTaylorO

OpenStudy (anonymous):

Thank you :)